OpenStudy (anonymous):

Please show your work and solve this b/6=18

OpenStudy (anonymous):

just multiply LHS and RHS by 6.

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

b/6 * 6 = 18 * 6

OpenStudy (anonymous):

now solve

OpenStudy (anonymous):

but whats the b? I know its any number but...

OpenStudy (anonymous):

b is any unknown value, you need to find

OpenStudy (anonymous):

what ans. you got?

OpenStudy (anonymous):

144 o.e

OpenStudy (anonymous):

Thats what the answer it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

in such type of problems you just have to perform suitable operations on both sides of equal.

OpenStudy (anonymous):

Oh OK, thanks! :)

OpenStudy (unklerhaukus):

\[\frac b6=18\]\[6\times\frac b6=6\times18\]\[\cancel6\times\frac b{\cancel 6}=6\times(10+8)\]\[b=(6\times10)+(6\times8)\]\[b=60+48\]\[b=\]